OpenStudy (anonymous):

What must be the nozzle velocity of the water from a fire hose if it is to reach a point 30m directly above the nozzle?

OpenStudy (cwrw238):

use one of the equations of motion for constant accelaeration in this case this is acceleration due to gravity 9.81 m s^-2 v^2 = u^2 + 2as u = initial velocity =? final velocity v = 0 acceleration = -9.81 distance s = 30 plug these values into above equation and solve for u

OpenStudy (cwrw238):

so you have 0 = u^2 - 2*9.81*30

OpenStudy (cwrw238):

so u^2 = 588.6 take the square root and you have it

OpenStudy (anonymous):

Thank you, makes sense